-2(y-4)=20-2y-12 How would this problem be solved?
Answer by stanbon(75887)

-2(y-4)=20-2y-12 How would this problem be solved?
---------------
-2y+8 = 20-2y-12
----
8 = 8
------
That means the equation is true for ALL values
of "y".
----
Cheers,
Stan H.
-------------